If a proton had a radius R and the charge was uniformly distributed, calculate using Bohr theory, the ground state energy of a H – atom when R = 10 Angstrom
 
 
 
 
 

Grade:12

2 Answers

Vikas TU
7 years ago
radius of atom => R =0.0529*n^2/z
and
Energy = 13.6*z^2/n^2
dividing both the eqns. we get,
r/E = 0.0529/13.6z
E = r*13.6*z/0.0529 => 10*10^-10*13.6*1/0.0529 => 2.57 * 10^-7 J
G Sir Askiitians faculty
6 years ago
For a point nucleus in H-atom:Ground state:2 2201 , – .4 B Bmv emvrr r πε = h =2 22 2 201 1 . 4πε ∴ = + B B   Bemm r r rh2024. rB 0.51Am eπε °∴ = = hPotential energy2 2 22 2 201 1 – . –27.2 ; . . 13.6eV4 B 2 2 B 2 Be mv eV K E mr r m r mr    = = = = = + π h hFor an spherical nucleus of radius R,If R > rB: the electron moves inside the sphere with radiusBr′ ( Br′ = new Bohr radius).Charge inside 343BBrreR ′  ′ =  